Grade 8 question
A rectangular lawn,4m by 4m, has a path around it that is the same width everywhere. The area of the path is 1.25 times as big as the area of the other lawn. What is the width of the path?

Answer by ankor@dixie-net.com(22740)

A rectangular lawn, 4m by 4m, has a path around it that is the same width everywhere.
The area of the path is 1.25 times as big as the area of the other lawn.
What is the width of the path?
:
4 * 4 = 16 sq/m the area of the garden
then
1.25(16) = 20 sq/m the area of the path
therefore
16 + 20 = 36 sq/m total area of the path and the garden
:
Let x = the width of the path, 2x is added to the overall dimensions
The area is square, therefore:
(2x+4)^2 = 36
Find the square root of both sides
2x + 4 = 6
2x = 6 - 4
2x = 2
x = 1 meter is the width of the path